# LHRS-Bot: Empowering Remote Sensing with VGI-Enhanced Large Multimodal Language Model

### v3

@@ Abstract
The revolutionary capabilities of large language models (LLMs) have paved the way for multimodal large language models (MLLMs) and fostered diverse applications across various specialized domains. In the remote sensing (RS) field, however, the diverse geographical landscapes and varied objects in RS imagery are not adequately considered in recent MLLM endeavors. To bridge this gap, we construct a large-scale RS image-text dataset, LHRS111LHRS stands for ’Language Helps Remote Sensing’.-Align, and an informative RS-specific instruction dataset, LHRS-Instruct, leveraging the extensive volunteered geographic information (VGI) and globally available RS images. Building on this foundation, we introduce LHRS-Bot, an MLLM tailored for RS image understanding through a novel multi-level vision-language alignment strategy and a curriculum learning method. Additionally, we introduce LHRS-Bench, a benchmark for thoroughly evaluating MLLMs’ abilities in RS image understanding. Comprehensive experiments demonstrate that LHRS-Bot exhibits a profound understanding of RS images and the ability to perform nuanced reasoning within the RS domain222Code and model weights are available at https://github.com/NJU-LHRS/LHRS-Bot.
@@
Figure 1: Overall performance comparison between LHRS-Bot and existing MLLMs on different RS image understanding tasks.
@@
Among various specialized domains, remote sensing (RS) image understanding is particularly important, as it directly involves the monitoring and analysis of the Earth’s surface and environment [72, 50, 64, 5]. However, RS images present significant challenges for holistic image understanding due to the intricate complexity of different objects across diverse landscapes worldwide, coupled with the variability in visual scales [44, 22, 51]. MLLMs bring together vision and language, offering possibilities for a unified and interactive understanding of RS images [64]. In this direction, several RS-specific MLLMs have been proposed, such as RSGPT [24], GeoChat [28], and SkyEyeGPT [78]. However, they are struggling to fit existing public RS datasets, due to the oversight of exploiting the extensive and worldwide RS features to infuse comprehensive RS visual knowledge into LLMs. Furthermore, these contemporary approaches of bridging vision and language domains tend to focus on the incorporation of high-level visual semantics. This strategy neglects the fact that different levels of visual information are crucial to fully align the linguistic and visual domains, which could facilitate a unified RS image understanding at different granularity [44, 67, 62].
To bridge the gap, we present LHRS-Bot, a specialized MLLM for RS, enhanced by globally available volunteered geographical information (VGI) and worldwide RS images. Specifically, we construct a large-scale, semantically rich, and feature-diverse dataset, LHRS-Align, by geographically pairing RS images from Google Earth333https://www.google.com/earth/ with plentiful attributed information from the OpenStreetMap VGI database444https://www.openstreesmap.org/, and consequently generating image captions using LLMs. Our rigorous data cleaning process, which includes deduplication, pruning, and semantic balancing, ensures that LHRS-Align comprises 1.15 million meaningful and high-quality RS image-text pairs, laying a solid foundation for embedding rich RS visual knowledge into MLLMs. Furthermore, we construct our instruction dataset, LHRS-Instruct, by reorganizing a broad range of open-source RS datasets into multi-task instructional datasets and employing GPT-4 [45] to create complex instructional data from the high-quality subset of LHRS-Align. Building on these datasets, we propose a novel RS-specific MLLM, LHRS-Bot, by implementing a novel bridging strategy for efficiently summarizing multi-level visual representations, alongside a curriculum learning methodology. LHRS-Bot achieves state-of-the-art performance across a variety of RS image understanding tasks (Fig. 1) and demonstrates remarkable abilities to detect intricate objects, engage with human conversation, and provide insights from visual information within RS images (Fig. 5).
@@ 2.1 Large-Scale RS Alignment Dataset——LHRS-Align
Data preparation. The RS images of LHRS-Align are sourced from Google Earth imagery, encompassing a worldwide coverage with a spatial resolution of 1 meter. The geographic features are gathered from OSM, a crowdsource VGI database containing abundant geographical objects. Each object (feature) in the OSM database is annotated as the digitization of the object geometry (e.g., polygons, lines, and points) as well as the semantic tags. A tag is a key-value pair that provides information about a geographic object555https://wiki.openstreetmap.org/wiki/Map_features, for example, “landuse=residential”. Thanks to the free access to the globally encompassing data sources, our pipeline can produce a dataset that offers a vast and comprehensive variety on a worldwide scale. For the current version of LHRS-Align, we collect RS images and OSM features within a scope of 0.3° longitude × 0.28° latitude for each of the selected cities from all over the world. We customize our own OSM database to streamline the subsequent feature selection process. More information about our data source and database is introduced in supplementary material.
@@
Caption generation with LLM. The use of image-caption pairs to incorporate visual knowledge into LLMs has been validated as an effective approach [33, 85, 73, 14, 3]. However, using plain image key-value labels can be problematic, as they lack grammatical structure, potentially disrupting the generative and comprehension abilities of LLMs. Similarly, rule-based caption generation strategies may introduce significant redundancy. Therefore, we opt to use open-source LLMs to generate image captions by summarising the key-value tags of the corresponding images. We experiment with various in-context learning examples and LLMs, ranging in size from 7B to 70B. After comprehensive testing by a domain expert, we select Vicuna-v1.5-13B [10] for caption generation as it achieves the best trade-offs in terms of generation efficiency and quality. Detailed information about the in-context examples and the specific deployment of Vicuna-v1.5-13B [10] can be found in supplementary material. Our final dataset, referred as LHRS-Align, comprises approximately 1.15 million image-caption pairs. Each pair is accompanied with additional metadata, such as image resolution and geographical location (country and city). Several examples of LHRS-Align are presented in supplementary material.
@@ 2.3 RS-specific MLLM Evaluation Benchmark——LHRS-Bench
We construct LHRS-Bench to evaluate the performance of MLLMs in the RS domain. LHRS-Bench employs hierarchical taxonomies to comprehensively evaluate MLLMs across various dimensions (Tab. 1). To prevent data leakage, we abstain from using public RS datasets and instead meticulously collect all the images from Google Earth. In terms of question format, we utilize single-choice questions [18, 40, 31, 30] with 2 to 4 candidate answers, enabling a more quantitative and objective evaluation compared to open-ended questions [4, 75, 6, 39]. Every question and answer within LHRS-Bench are manually crafted to guarantee precision and reliability. Totally, LHRS-Bench comprises 108 images and 690 questions. Some examples of LHRS-Bench are shown in supplementary material.
@@ 4.3 Evaluation on LHRS-Bench
We evaluate different MLLMs on our LHRS-Bench benchmark. For each question within the benchmark, choices are shuffled, and the generation process is repeated four times. An answer is considered incorrect if any attempt yields an incorrect response. Same prompts are used for evaluation, as detailed in supplementary material. We consider the model answers correctly if it outputs the correct choice or the answer corresponding to the correct choice. We calculate accuracy as our evaluation metric.
The results, presented in Tab. 5, show that LHRS-Bot achieves the best results on this benchmark. It is worth noting that, LHRS-Bot outperforms others in terms of the evaluation dimensions of spatial resolution and image modality, even though these aspects were not explicitly trained for in the models, highlighting the effectiveness of RS-specific MLLM training. However, although LHRS-Bot has achieved the highest performance, there is still significant room for improvement in MLLM capabilities within the RS domain. Examples of evaluation results are available on supplementary material.
@@ 4.5 Ablation Analysis
5 Related Work
MLLMs, specifically vision-language MLLMs, have extended the abilities of LLMs by incorporating visual alignment and visual instruction tuning [1, 34, 17, 33, 85, 3], enabling them to tackle cross-modal tasks in various domains [81, 71, 66, 74, 61, 58, 54, 55, 36, 15, 33, 43]. Concurrently, advancements in RS vision-language models, such as RSGPT [24], GeoChat [28], and SkyEyeGPT [78], have achieved promising improvements in understanding RS images. However, the present RS-specific MLLMs, fine-tuned using limited public task-specific datasets, neglect the potential of leveraging global RS images and vast crowdsourced databases to enhance LLMs with extensive RS visual knowledge. In contrast, we unleash the full potential of LLMs for RS applications by leveraging the abundant geographical information with global available RS images and employing a more effective multi-level alignment strategy. Further details about recent related development are available in supplementary material.
@@ 6 Conclusion
We propose LHRS-Bot, an MLLM for the RS domain. In our effort to unleash the full potential of LLM for RS image understanding, we curate a specialized large-scale alignment dataset, LHRS-Align, designed for RS-specific alignment, along with LHRS-Instruct, a RS multimodal instruction-following dataset, to facilitate the instruction following ability of LHRS-Bot. Furthermore, we establish a benchmark, LHRS-Bench, which provides a systematic evaluation framework for RS MLLMs. In comparative analyses, LHRS-Bot stands out for its exceptional RS domain knowledge, significantly outperforming both the general-purpose and RS-specific MLLMs.
While LHRS-Bot has shown strong capabilities in RS applications, we must acknowledge that it shares some limitations inherent to LLMs, such as susceptibility to producing hallucinations. Users are therefore advised to employ LHRS with caution. Concurrently, we believe that the development of higher quality alignment datasets and instructional datasets, along with improved training strategies, has the potential to significantly improve the performance of MLLMs in the interpretation of RS images.
Appendix 0.A offers an in-depth review of related work.
Appendix 0.B presents an explanation of the methods used to construct the LHRS-Align dataset.
Appendix 0.C outlines the steps taken to develop the LHRS-Instructinstruction dataset.
@@
Appendix 0.E provides detailed information about the evaluation test dataset.
@@ Appendix 0.B Details about LHRS-Align Dataset
0.B.1 Data Source.
For now, we have collected images that cover the 10000 most populous cities from certain countries. Due to challenges such as the scarcity of OSM data for certain cities (particularly in underdeveloped regions), the remaining images cover 9,259 cities across 129 countries, as shown in Fig. 6. Considering the spatial resolution limitation of GE images, we only utilize the polygon features of OSM for the current version of our dataset, with the extension of polyline and point features left to our future works. Furthermore, we discard the polygon features tagged with the keys “boundary” and “barrier”, as they do not exactly describe the function of the feature.
0.B.2 Geo-alignment and Image Processing.
We search the database and find out qualified features as anchors to locate the corresponding RS images. The principle is choosing features whose areas are larger than a 128×128 pixel image and have aspect ratios less than 4. This is to ensure that the chosen features are adequately large and do not exhibit excessive slenderness. For the RS images, we resize any image larger than 768×768 pixels down to 768×768. Subsequently, the images undergo a deduplication process and a pruning step, which involves removing images that predominantly feature vast ocean areas or are obscured by clouds. The latter process is facilitated by employing a trained classification network. These postprocessing procedures reduce the samples from 7B to 4B. Based on the geographical extent of each image, the corresponding features are selected out whose spatial area is smaller than 1/64 of the image’s size. The spatial areas of the features and images are calculated in Web Mercator projection.
0.B.3 Filtering Keys.
@@
0.B.5 Details about Generating Captions.
@@ 0.C.1 Instruction-Following Dataset from Public RS Caption Dataset
Image deduplication is conducted using the fastdup tool777https://github.com/visual-layer/fastdup?tab=readme-ov-file.

### v4

@@ Abstract
The revolutionary capabilities of large language models (LLMs) have paved the way for multimodal large language models (MLLMs) and fostered diverse applications across various specialized domains. In the remote sensing (RS) field, however, the diverse geographical landscapes and varied objects in RS imagery are not adequately considered in recent MLLM endeavors. To bridge this gap, we construct a large-scale RS image-text dataset, LHRS111LHRS stands for ’Language Helps Remote Sensing’.-Align, and an informative RS-specific instruction dataset, LHRS-Instruct, leveraging the extensive volunteered geographic information (VGI) and globally available RS images. Building on this foundation, we introduce LHRS-Bot, an MLLM tailored for RS image understanding through a novel multi-level vision-language alignment strategy and a curriculum learning method. Additionally, we introduce LHRS-Bench, a benchmark for thoroughly evaluating MLLMs’ abilities in RS image understanding. Comprehensive experiments demonstrate that LHRS-Bot exhibits a profound understanding of RS images and the ability to perform nuanced reasoning within the RS domain222Data, Code and model are available at https://github.com/NJU-LHRS/LHRS-Bot.
@@
Among various specialized domains, remote sensing (RS) image understanding is particularly important, as it directly involves the monitoring and analysis of the Earth’s surface and environment [72, 50, 64, 5]. However, RS images present significant challenges for holistic image understanding due to the intricate complexity of different objects across diverse landscapes worldwide, coupled with the variability in visual scales [44, 22, 51]. MLLMs bring together vision and language, offering possibilities for a unified and interactive understanding of RS images [64]. In this direction, several RS-specific MLLMs have been proposed, such as RSGPT [24], GeoChat [28], and SkyEyeGPT [78]. However, they are struggling to fit existing public RS datasets, due to the oversight of exploiting the extensive and worldwide RS features to infuse comprehensive RS visual knowledge into LLMs. Furthermore, these contemporary approaches of bridging vision and language domains tend to focus on the incorporation of high-level visual semantics. This strategy neglects the fact that different levels of visual information are crucial to fully align the linguistic and visual domains, which could facilitate a unified RS image understanding at different granularity [44, 67, 62]. Detailed related works are available in supplementary material.
To bridge the gap, we present LHRS-Bot, a specialized MLLM for RS, enhanced by globally available volunteered geographical information (VGI) and worldwide RS images. Specifically, we construct a large-scale, semantically rich, and feature-diverse dataset, LHRS-Align, by geographically pairing RS images with plentiful attributed information from the OpenStreetMap VGI database333https://www.openstreetmap.org/, and consequently generating image captions using LLMs. Our rigorous data cleaning process, which includes deduplication, pruning, and semantic balancing, ensures that LHRS-Align comprises 1.15 million meaningful and high-quality RS image-text pairs, laying a solid foundation for embedding rich RS visual knowledge into MLLMs. Furthermore, we construct our instruction dataset, LHRS-Instruct, by reorganizing a broad range of open-source RS datasets into multi-task instructional datasets and employing GPT-4 [45] to create complex instructional data from the high-quality subset of LHRS-Align. Building on these datasets, we propose a novel RS-specific MLLM, LHRS-Bot, by implementing a novel bridging strategy for efficiently summarizing multi-level visual representations, alongside a curriculum learning methodology. LHRS-Bot achieves state-of-the-art performance across a variety of RS image understanding tasks (Fig. 1) and demonstrates remarkable abilities to detect intricate objects, engage with human conversation, and provide insights from visual information within RS images (Fig. 5).
@@ 2.1 Large-Scale RS Alignment Dataset——LHRS-Align
Data Preparation. The RS images of LHRS-Align are sourced from Google Earth imagery444https://www.google.com/earth/, encompassing a worldwide coverage with a spatial resolution of 1 meter. It should be noted that our proposed pipeline is applicable to any RS imagery with geographic coordinates. The geographic features are gathered from OSM, a crowdsource VGI database containing abundant geographical objects. Each object (feature) in the OSM database is annotated as the digitization of the object geometry (e.g., polygons, lines, and points) as well as the semantic tags. A tag is a key-value pair that provides information about a geographic object555https://wiki.openstreetmap.org/wiki/Map_features, for example, “landuse=residential”. Thanks to the free access to the globally encompassing data sources, our pipeline can produce a dataset that offers a vast and comprehensive variety on a worldwide scale. For the current version of LHRS-Align, we collect RS images and OSM features within a scope of 0.3° longitude × 0.28° latitude for each of the selected cities from all over the world. We customize our own OSM database to streamline the subsequent feature selection process. More information about our data source and database is introduced in supplementary material.
@@
Caption Generation with LLM. The use of image-caption pairs to incorporate visual knowledge into LLMs has been validated as an effective approach [33, 85, 73, 14, 3]. However, using plain image key-value labels can be problematic, as they lack grammatical structure, potentially disrupting the generative and comprehension abilities of LLMs. Similarly, rule-based caption generation strategies may introduce significant redundancy. Therefore, we opt to use open-source LLMs to generate image captions by summarising the key-value tags of the corresponding images. We experiment with various in-context learning examples and LLMs, ranging in size from 7B to 70B. After comprehensive testing by a domain expert, we select Vicuna-v1.5-13B [10] for caption generation as it achieves the best trade-offs in terms of generation efficiency and quality. Detailed information about the in-context examples and the specific deployment of Vicuna-v1.5-13B [10] can be found in supplementary material. Our final dataset, referred as LHRS-Align, comprises approximately 1.15 million image-caption pairs. Each pair is accompanied with additional metadata, such as image resolution and geographical location (country and city). Several examples of LHRS-Align are also presented in supplementary material.
@@ 2.3 RS-specific MLLM Evaluation Benchmark——LHRS-Bench
We construct LHRS-Bench to evaluate the performance of MLLMs in the RS domain. LHRS-Bench employs hierarchical taxonomies to comprehensively evaluate MLLMs across various dimensions (Tab. 1). To prevent data leakage, we abstain from using public RS datasets and instead meticulously collect images from Google Earth.In terms of question format, we utilize single-choice questions
| Dimension | Sub-dimension | Count |
| Recognition | Identity | 634 |
| Color | 113 | |
| Orientation | 39 | |
| Shape | 37 | |
| Area | 75 | |
| Imagery | Resolution | 21 |
| Modality | 23 | |
| Spatial-awareness | Location | 204 |
| Distance | 22 | |
| Quantity | - | 137 |
| Reasoning | - | 46 |
[18, 40, 31, 30] with 2 to 4 candidate answers, enabling a more quantitative and objective evaluation compared to open-ended questions [4, 75, 6, 39]. Every question and answer within LHRS-Bench are manually crafted to guarantee precision and reliability. Totally, LHRS-Bench comprises 108 images and 690 questions. Some examples of LHRS-Bench are shown in supplementary material.
@@ 4.2 Quantitative Evaluation
| Method | RSVG | DIOR-RSVG | Avg. |
| QWen-VL-Chat | 44.76 | 80.65 | 62.71 |
| MiniGPTv2 | 46.64 | 85.99 | 66.32 |
| SkyEyeGPT | 70.50 | 88.59 | 79.55 |
| LHRS-Bot | 73.45 | 88.10 | 80.78 |
@@ 4.3 Evaluation on LHRS-Bench
We evaluate LHRS-Bot with different open-source and closed-source MLLMs on our LHRS-Bench benchmark. For each question within the benchmark, choices are shuffled, and the generation process is repeated four times. An answer is considered incorrect if any attempt yields an incorrect response. Same prompts are used for evaluation, as detailed in supplementary material. We consider the model answers correctly if it outputs the correct choice or the answer corresponding to the correct choice. We calculate accuracy as our evaluation metric.
The results presented in Tab. 5 indicate that LHRS-Bot outperforms all other open-source models. It is worth noting that, LHRS-Bot greatly outperforms other open-source models in terms of the evaluation dimensions of spatial resolution and image modality, even though these aspects were not explicitly trained for in the models, highlighting the effectiveness of RS-specific MLLM training. Additionally, LHRS-Bot outperforms the strong closed-source model Claude and even matches the performance of GPT-4V. However, there is still significant room for improvement in MLLM capabilities within the RS domain. Examples of evaluation results are available on supplementary material.
@@ 5 Conclusion
We propose LHRS-Bot, an MLLM for the RS domain. To unleash the potential of LLMs for RS image understanding, we curate a large-scale dataset, LHRS-Align, for RS-specific alignment, and LHRS-Instruct, a multimodal instruction-following dataset to enhance LHRS-Bot’s instruction-following capabilities. Additionally, we introduce LHRS-Bench, a benchmark providing a systematic evaluation framework for RS MLLMs.
In comprehensive comparison experiments, LHRS-Bot demonstrates exceptional RS domain knowledge, surpassing both general-purpose and RS-specific MLLMs. Despite its impressive performance, LHRS-Bot shares limitations common to LLMs, such as susceptibility to hallucinations. We believe that developing higher-quality alignment and instructional datasets, along with improved training strategies, can further enhance the performance of MLLMs for interpreting RS images.
Acknowledgement
This research was supported by the National Natural Science Foundation of China under Grant 42071297, and in part by the AI & AI for Science Project of Nanjing University under Grant 02091480605203. We are grateful to High Performance Computing Center of Nanjing University for their help on GPU resources. We also would like to thank the anonymous reviewers for their constructive comments.
Appendix 0.A givens an in-depth review of related work.
Appendix 0.B presents explanation of the methods used to construct the LHRS-Align dataset.
Appendix 0.C outlines the steps taken to develop the LHRS-Instruct dataset.
@@ Acknowledgement
Appendix 0.E provides detailed information about the evaluation datasets.
@@ Appendix 0.B Details about LHRS-Align Dataset
0.B.1 Data Source
We have collected images that cover the most populous cities from certain countries. Due to challenges such as the scarcity of OSM data for certain cities (particularly in underdeveloped regions), the remaining images cover 9,259 cities across 129 countries. Considering the spatial resolution limitation of RS images, we only utilize the polygon features of OSM for the current version of our dataset, with the extension of polyline and point features left to our future works. Furthermore, we discard the polygon features tagged with the keys “boundary” and “barrier”, as they do not exactly describe the function of the feature.
0.B.2 Geo-alignment and Image Processing
We search the database and find out qualified features as anchors to locate the corresponding RS images. The principle is choosing features whose areas are larger than a 128×128 pixel image and have aspect ratios less than 4. This is to ensure that the chosen features are adequately large and do not exhibit excessive slenderness. For the RS images, we resize any image larger than 768×768 pixels down to 768×768. Subsequently, the images undergo a deduplication process with fastdup 777https://github.com/visual-layer/fastdup?tab=readme-ov-file and a pruning step, which involves removing images that predominantly feature vast ocean areas or are obscured by clouds. The latter process is facilitated by employing a trained classification network. These postprocessing procedures reduce the samples from 7M to 4M. Based on the geographical extent of each image, the corresponding features are selected out whose spatial area is smaller than 1/64 of the image’s size. The spatial areas of the features and images are calculated in Web Mercator projection.
0.B.3 Filtering Keys
@@ 0.B.4 Semantic Balancing
0.B.5 Details about Generating Captions
@@ 0.C.1 Instruction-Following Dataset from Public RS Caption Dataset
Image deduplication is conducted using the fastdup tool.